# Cobordism Obstructions to Complex Sections II: Torsion Obstructions

### v1

@@ Abstract.
In the previous paper, we studied the existence of complex sections on almost complex manifolds up to cobordism. We proved a necessary and sufficient condition for their existence in $\Omega_{*}^{U}\otimes\mathbb{Q}$ in terms of the Chern classes. In this paper we study the existence question in $\Omega^{U}_{*}$. We call corresponding obstructions torsion and we determine them for small values of $r$. Calculations with the Adams-Novikov spectral sequence for the Thom spectra defined in the previous paper allow us to show the vanishing of the torsion obstruction for low $r$. For prime $p$, we show the vanishing of $p$ torsion obstructions for finding $r$ complex sections for $r<p^{2}-p$.
@@ 1. Introduction
There is a classical problem to determine whether a manifold admits $r$ linearly independent tangent vector fields. In the work of Bokstedt, Dupont and Svane [4] and in the previous paper [9], we approached this problem by instead determining the obstruction to finding a cobordant manifold with $r$ vector fields. We were able to describe the rational obstruction to $r$ linearly independent complex sections of the tangent bundle of an almost complex manifold up to cobordism. This obstruction is given in terms of Chern characteristic numbers as follows:
@@
Let $M^{2d}$ be a $d$-dimensional almost-complex manifold. The Chern characteristic classes $s_{\omega}(M^{2d})=0$ for all $\omega$ of length greater than or equal to $d-r+1$ if and only if there exists integer $c$ such that the cobordism class $c[M^{2d}]$ contains a manifold $N^{2d}$ with $r$ complex sections on $TN$.
For any cobordism class $[M^{2d}]$ with vanishing rational obstruction, we can ask whether there are any divisibility conditions for the constant $c$ in the above theorem. If, for all $[M^{2d}]\in\Omega_{2d}^{U}$ with vanishing rational obstruction, the constant $c$ may be chosen to be $1$ mod $p$ for prime $p$, we say that the $p$ torsion obstruction vanishes. This is equivalent to a particular map having no $p$-torsion in its image, see Theorem 4.6 in [9]. We will show in this paper, Proposition 3.6, that any $p$-torsion in the image of this map is equivalent to a non-vanishing differential in the $p$-primary Adams-Novikov spectral sequence for $\mathbf{MTU}(d-r)$. The prime $p>2$ case is tackled first:
Theorem 1.2.
For $p>2$, the $p$-torsion obstruction to the existence of $r$ complex sections vanishes when $r<p^{2}-p$.
@@
We can make a further improvement to the above result when $d$ is odd.
Theorem 1.4.
@@ 2.1. The spectra MTU(d) and MTU(d,r)
Here we introduce the spectra $\mathbf{MTU}(d)$ and $\mathbf{MTU}(d,r)$. The homotopy groups of these spectra will be the natural objects of study. The real case is described in [4] and [7]. Most of this work is repeated from our previous paper [9]. It should be noted that there is some disagreement about the proper indexing of these spectra in [4] and [7]. We follow the convention given in Bokstedt, Dupont and Svane. [4] Write $G_{\mathbb{C}}(d,n)$ for the complex Grassmannian of $d$ dimensional complex subspaces of $\mathbb{C}^{d+n}$. Let $U_{\mathbb{C},d,n}\rightarrow G_{\mathbb{C}}(d,n)$ be the tautological $d$ dimensional complex vector bundle and let $U_{\mathbb{C},d,n}^{\perp}\rightarrow G_{\mathbb{C}}(d,n)$ be the $n$ dimensional orthogonal complement of $U_{\mathbb{C},d,n}$.
@@
There exists a canonical map $G_{\mathbb{C}}(d,n)\rightarrow G_{\mathbb{C}}(d,n+1)$ defined using the composition:
@@
There is also a map $G(d-r,n)\rightarrow G(d,n)$ which takes a $(d-r)$-complex plane $P\subset\mathbb{C}^{d-r+n}$ to the $d$-plane $P\oplus\mathbb{C}^{r}\subset\mathbb{C}^{d-r+n}\oplus\mathbb{C}^{r}$. Under this map, the pullback of $U_{\mathbb{C},d,n}^{\perp}$ is $U_{\mathbb{C},d-r,n}^{\perp}$. Thus, there are maps of Thom spaces $\mathbf{MTU}(d-r)_{2n}\rightarrow\mathbf{MTU}(d)_{2n}$. Since this map commutes with the spectrum map, it defines a map of spectra $\mathbf{MTU}(d-r)\rightarrow\mathbf{MTU}(d)$.
@@
The following commutative diagram relates these two constructions and shows that they are the same.
@@
$\pi_{2d}(\mathbf{MTU}(d-r))$ is the even complex section cobordism group of $2d$ dimensional manifolds with $r$ complex sections. That is, elements of $\pi_{2d}(\mathbf{MTU}(d-r))$ are classes of almost complex manifolds with $r$ complex sections. Two classes are the same if there is a complex cobordism equipped with $r+1$ sections between them. The map $\pi_{2d}(\mathbf{MTU}(d-r))\rightarrow\pi_{2d}(\mathbf{MU})$ is the forgetful map. The image of the forgetful map is all cobordism classes which contain a manifold which can be equipped with $r$ complex sections.
The space $\mathbf{MTU}(d,1)$ can be given a simpler description which will be useful for computing the homotopy groups:
@@
By the second construction, we observe that $\mathbf{MTU}(d,1)_{2n}\cong Th(p^{*}_{W_{\mathbb{C},1}}U^{\perp}_{\mathbb{C},d,n})/Th(p^{*}_{V_{\mathbb{C},1}}U^{\perp}_{\mathbb{C},d,n})$ where $p^{*}_{V_{\mathbb{C},1}}:V_{1}(U_{\mathbb{C},d,n})\cong S(U_{\mathbb{C},d,n})\rightarrow G_{\mathbb{C}}(d,n)$ and $p^{*}_{V_{\mathbb{C},1}}:W_{1}(U_{\mathbb{C},d,n})\cong D(U_{\mathbb{C},d,n})\rightarrow G_{\mathbb{C}}(d,n)$.
There is a property that for two vector bundles $V_{1}$ and $V_{2}$ over the same base,
@@
$\mathbf{MTU}(d,1)_{2n}\cong Th(U_{\mathbb{C},d,n}\oplus U^{\perp}_{\mathbb{C},d,n})\cong Th(\mathbb{R}^{2d+2n}\times G_{\mathbb{C}}(d,n))\cong S^{2d+2n}\vee\Sigma^{2d+2n}G_{\mathbb{C}}(d,n)$
Letting $n\rightarrow\infty$ proves the result. ∎
@@
This result follows by dualizing the corresponding result Proposition 2.2 in cohomology from [9]. ∎
@@
There is an obstruction to lifting an element of $\pi_{2d}(\mathbf{MU})$ to $\pi_{2d}(\mathbf{MTU}(d-r))$ which lands in the homotopy group $\pi_{2d}(\mathbf{MTU}(d,r))$. In [9] we described the image of $\pi_{*}(\mathbf{MTU}(d))\otimes\mathbb{Q}\rightarrow\pi_{*}(MU)\otimes\mathbb{Q}$ as follows:
@@
Let $M^{2d}$ be a $d$-dimensional almost-complex manifold. The Chern characteristic classes $s_{\omega}(M^{2d})=0$ for all $\omega$ of length greater than or equal to $d-r+1$ if and only if the cobordism class $[M^{2d}]$ lifts to $\pi_{2d}(\mathbf{MTU}(d-r))\otimes\mathbb{Q}$. Geometrically, this implies there exists integer $c$ such that the cobordism class $c[M^{2d}]$ contains a manifold $N^{2d}$ with $r$ complex sections on $TN$.
In this document, we begin to describe the image of the forgetful map $\pi_{*}(\mathbf{MTU}(d))\rightarrow\pi_{*}(\mathbf{MU})$. As stated in the introduction, we will call the torsion in the cokernel, the torsion obstructions. Thus we determine which cobordism classes of manifolds can be equipped with complex sections. To do so, we will study the $p$-local Adams-Novikov spectral sequence for $\mathbf{MTU}(d)$.
@@ 3. Computing the torsion obstruction
This section will be devoted to identifying which elements of $\pi_{*}(\mathbf{MU})$ can be lifted to $\pi_{*}(\mathbf{MTU}(d))$. We will use the following notational conventions. Then let $B_{i}\in MU_{*}(\mathbf{MU})=\pi_{*}(\mathbf{MU})\otimes H_{*}(\mathbf{MU})$, be the generators (as a $\pi_{*}(MU)$ algebra). These $B_{i}$ correspond to the elements $b_{i}\in H_{*}(\mathbf{MU})\cong\mathbb{Z}[b_{1},b_{2},...]$. There is an injective Hurewicz homomorphism $\pi_{*}(\mathbf{MU})\rightarrow H_{*}(\mathbf{MU})$. By slight abuse of notation, we call the elements of $\pi_{*}(\mathbf{MU})$ by their image in $H_{*}(\mathbf{MU})$. We begin by stating the following results about $MU$ theory originally developed by [10] and described in detail in [11] Appendix 2.1.
@@
  • A right unit corresponding to the Hurewicz homomorphism $\eta_{R}:\pi_{*}(\mathbf{MU})\rightarrow MU_{*}(\mathbf{MU})$, defined by the formula:
@@
For each prime $p$, there is an associative ring spectrum $\mathbf{BP}$ which is a retract of $\mathbf{MU}_{(p)}$, moreover, $\mathbf{MU}_{(p)}$ splits as copies of $\mathbf{BP}$. The homotopy group $\pi_{*}(\mathbf{BP})$ is $\mathbb{Z}_{(p)}[v_{1},v_{2},...]$ where $v_{i}\in\pi_{2(p^{i}-1)}(\mathbf{BP})$. The pair $(\pi_{*}(\mathbf{BP}),BP_{*}(\mathbf{BP}))$ is a Hopf algebroid. We will write $BP_{*}(\mathbf{BP})=\pi_{*}(\mathbf{BP})[t_{1},t_{2},...]$ with $t_{i}$ in degree $2(p^{i}-1)$. The algebroid has the structure:
@@
  • A right unit corresponding to the Hurewicz homomorphism $\eta_{R}:\pi_{*}(\mathbf{BP})\rightarrow BP_{*}(\mathbf{BP})$, defined by the formula:
@@
The basis $m_{i}$ comes from the elements $[\mathbb{C}P^{p^{i}-1}]/p$. The generators $v_{i}$ are defined by
@@
Let $X$ be a spectrum with torsion-free homology concentrated in even degrees. Let $E$ be either $\mathbf{MU}$ or $\mathbf{BP}$ for some prime $p$. Then, $E_{*}(X)\cong\pi_{*}(E)\otimes_{\mathbb{Z}}E_{*}(X;\mathbb{Z})$ and is torsion-free concentrated in even degrees.
@@
Let $E$ be either $\mathbf{MU}$ or $\mathbf{BP}$ for some prime $p$. The map
$E_{*}(\mathbf{MTU}(d))\rightarrow E_{*}(\mathbf{MU})=\pi_{*}(E)[B_{1},B_{2},...]$
is an injection, which is an isomorphism in degrees less than $22d+$ 2. Moreover, both modules are concentrated in even degrees. The image of $E_{*}(\mathbf{MTU}(d))$ maps onto polynomials of degree $d$.
$E_{*}(\mathbf{MU})\rightarrow E_{*}(\overline{\mathbf{MTU}}(d))$
is the quotient map killing all polynomials of degree $d$ or less. These maps induce natural structures as $E_{*}(E)$ comodules.
We can now describe the $p$-local Adams-Novikov spectral sequence, see [10] or [11].
@@
For an element in $E_{q}^{s,t}$, we will refer to $s$ as the resolution of the element and $t$ as the degree. The element, if it survives, corresponds to an element of $\pi_{t-s}$. Consider the Adams-Novikov spectral sequence for $\mathbf{MU}_{(p)}$. Since $\mathbf{MU}_{(p)}$ splits as a copy of $\mathbf{BP}$’s, $E_{2}^{s,*}(\mathbf{MU}_{(p)})=0$ for $s>0$. So the spectral sequence collapses, and $E_{2}^{0,*}(\mathbf{MU}_{(p)})=\pi_{*}(\mathbf{MU}_{(p)})$. In particular, it is a free $\mathbb{Z}_{(p)}$ module. Moreover,
$E_{2}^{0,*}(\mathbf{MTU}(d)_{(p)})=\mathbf{Hom}(BP_{*}(\mathbf{MTU}(d)_{(p)}),\pi_{*}(\mathbf{BP}))$
We are primarily concerned with the image of $BP_{*}(\mathbf{MTU}(d)_{(p)})$, so we only need to know the image of the map $E_{2}^{0,*}(\mathbf{MTU}(d))\rightarrow E_{2}^{0,*}(\mathbf{MU})$. This is entirely described by the map
$\mathbf{Hom}_{BP_{*}(\mathbf{BP})}(\pi_{*}(\mathbf{BP}),BP_{*}(\mathbf{MTU}(d)_{(p)}))\rightarrow\mathbf{Hom}_{BP_{*}(\mathbf{BP})}(\pi_{*}(\mathbf{BP}),BP_{*}(\mathbf{MU}_{(p)}))=\pi_{*}(\mathbf{MU}_{(p)})$
This map has cokernel which is a free $\mathbb{Z}_{(p)}$ module. The following proposition follows:
@@
Moreover, $[M]$ lifts to $\pi_{*}(\mathbf{MTU}(d))$ if and only if there is no Adams differential originating from $x$. Moreover, any free part of $\pi_{*}(\mathbf{MTU}(d))$ comes from the zero line.
Thus, the differentials in the Adams-Novikov spectral sequence for $\mathbf{MTU}(d)$ correspond with torsion obstructions. We will use the cobar complex in order to compute the $E_{2}$ page. See Appendix 1 in [11] for more details. Recall that $BP_{*}(\mathbf{BP})$ has an action of $\pi_{*}(\mathbf{BP})$ given by multiplication by the image of $\eta_{R}$. Let $\overline{\Gamma}$ be the quotient of $BP_{*}(\mathbf{BP})$ the image of $\eta_{L}$.
@@
Let $(M,\psi)$ be a left $BP_{*}(\mathbf{BP})$ comodule. Define the cobar complex as the tensor product $C_{BP}^{s}(M)=\overline{\Gamma}^{\otimes s}\otimes_{\pi_{*}(\mathbf{BP})}M$ The differential $d_{1}:C_{BP}^{s}(M)\rightarrow C_{BP}^{s+1}(M)$ is given by:
@@
It will typically be convenient to view the cobar complex as the $E_{1}$ page of the Adams-Novikov spectral sequence. We observe that the 00 line of the $E_{2}$ page of the Adams-Novikov spectral sequence is generated by primitives, i.e. elements of the form $\psi(x)=1\otimes x$ where $\psi$ is the coaction. The following proposition will be used to compute $E_{2}(\mathbf{MTU}(d))$

### v2

@@ Abstract.
In the previous paper, we studied obstructions to the existence of complex sections on almost complex manifolds up to cobordism. We determined the obstruction rationally, in terms of the Chern classes. In this paper, we study the torsion obstructions, that is, the obstructions which vanish after tensoring with $\mathbb{Q}$ or multiplication by an integer. Calculations with the Adams-Novikov spectral sequence for the Thom spectra $\mathbf{MTU}(d)$ allow us to show the torsion obstructions for low $r$. For prime $p\geq 3$, we show that torsion obstructions for finding $r$ complex sections of order $p$ vanish for $r<p^{2}-p$.
Key words and phrases:
complex cobordism, complex sections, spectra, cobordism category.
@@ 1. Introduction
There is a classical problem to determine whether a manifold admits $r$ linearly independent tangent vector fields. In the work of Bökstedt, Dupont and Svane [4] and in the previous paper [9], we approached this problem by instead determining the obstruction to finding a cobordant manifold with $r$ vector fields. We were able to describe the rational obstruction to $r$ linearly independent complex sections of the tangent bundle of an almost complex manifold up to cobordism. This obstruction is given in terms of Chern characteristic numbers as follows:
@@
Let $M^{2d}$ be a $d$-dimensional almost-complex manifold. The Chern characteristic classes $s_{\omega}(M^{2d})=0$ for all partitions $\omega$ of length greater than $d-r$ if and only if there exists integer $c$ such that the cobordism class $c[M^{2d}]$ contains a manifold $N^{2d}$ with $r$ complex sections on $TN$.
The length of a partition is the number of summands in the partition. For any cobordism class $[M^{2d}]$ with vanishing rational obstruction, we can ask whether there are any divisibility conditions for the constant $c$ in the above theorem. If, for all $[M^{2d}]\in\Omega_{2d}^{U}$ with vanishing rational obstruction, the constant $c$ may be chosen to be $1$ mod $p$ for prime $p$, we say that the $p$ torsion obstruction vanishes. This is equivalent to a particular map having no $p$-torsion in its image, see Theorem 4.6 in [9]. We will show in this paper, Proposition 3.6, that $p$ torsion obstructions correspond to non-vanishing differentials in the $p$-primary Adams-Novikov spectral sequence for $\mathbf{MTU}(d-r)$. The prime $p>2$ case is tackled first:
Theorem 1.2 (Theorem 3.9).
For $p>2$, the $p$-torsion obstruction to the existence of $r$ complex sections of the tangent bundle of an almost complex manifold vanishes when $r<p^{2}-p$.
@@
In the case $d$ is odd, we can go one step further:
Theorem 1.4 (Theorem 3.12).
@@
Acknowledgments: I would like to thank all the people who have helped me as I have completed this paper. First, I would like to thank the University of Oregon Department of Mathematics for supporting me during this work. I most especially thank my advisor, Boris Botvinnik, for the time he spent guiding and advising me throughout this project. I also thank Douglas Ravenel for his feedback and suggestions.
@@ 2.1. The spectra MTU(d) and MTU(d,r)
Here we introduce the spectra $\mathbf{MTU}(d)$ and $\mathbf{MTU}(d,r)$. The homotopy groups of these spectra will be the natural objects of study. The real case is described in [4] and [7]. Most of this work is repeated from our previous paper [9]. It should be noted that there is some disagreement about the proper indexing of these spectra between [4] and [7]; we follow the convention given in Bökstedt, Dupont and Svane. [4] Write $G_{\mathbb{C}}(d,n)$ for the complex Grassmannian of $d$ dimensional complex subspaces of $\mathbb{C}^{d+n}$. Let $U_{\mathbb{C},d,n}\rightarrow G_{\mathbb{C}}(d,n)$ be the tautological $d$ dimensional complex vector bundle and let $U_{\mathbb{C},d,n}^{\perp}\rightarrow G_{\mathbb{C}}(d,n)$ be the $n$ dimensional orthogonal complement of $U_{\mathbb{C},d,n}$.
@@
There exists a canonical map $G_{\mathbb{C}}(d,n)\rightarrow G_{\mathbb{C}}(d,n+1)$ defined by the composition:
@@
There is also a map $G(d-r,n)\rightarrow G(d,n)$ which takes a $(d-r)$-dimensional complex plane $P\subset\mathbb{C}^{d-r+n}$ to the $d$-plane $P\oplus\mathbb{C}^{r}\subset\mathbb{C}^{d-r+n}\oplus\mathbb{C}^{r}$. Under this map, the pullback of $U_{\mathbb{C},d,n}^{\perp}$ is $U_{\mathbb{C},d-r,n}^{\perp}$. The pullback induces canonical maps of Thom spaces $\mathbf{MTU}(d-r)_{2n}\rightarrow\mathbf{MTU}(d)_{2n}$. Since this map commutes with the spectrum map, it defines a map of spectra $\mathbf{MTU}(d-r)\rightarrow\mathbf{MTU}(d)$.
@@
The following commutative diagram relates these two constructions and shows that they are homotopy equivalent.
@@
$\pi_{2d}(\mathbf{MTU}(d-r))$ is the even complex section cobordism group of $2d$ dimensional manifolds with $r$ complex sections. Elements of $\pi_{2d}(\mathbf{MTU}(d-r))$ are classes of almost complex manifolds with $r$ complex sections. Two classes are the same if there is a complex cobordism equipped with $r+1$ sections between them. The map $\pi_{2d}(\mathbf{MTU}(d-r))\rightarrow\pi_{2d}(\mathbf{MU})$ is the forgetful map. The image of the forgetful map is all cobordism classes which contain a manifold which can be equipped with $r$ complex sections.
The space $\mathbf{MTU}(d,1)$ can be given a simpler description which will be useful for computing the homotopy groups: (cf. [4][Proposition 2.6])
@@
We observe that $\mathbf{MTU}(d,1)_{2n}\cong Th(p^{*}_{W_{\mathbb{C},1}}U^{\perp}_{\mathbb{C},d,n})/Th(p^{*}_{V_{\mathbb{C},1}}U^{\perp}_{\mathbb{C},d,n})$ where $p^{*}_{V_{\mathbb{C},1}}$ is the projection $p^{*}_{V_{\mathbb{C},1}}:V_{1}(U_{\mathbb{C},d,n})\cong S(U_{\mathbb{C},d,n})\rightarrow G_{\mathbb{C}}(d,n)$ and $p^{*}_{W_{\mathbb{C},1}}$ is the projection $p^{*}_{W_{\mathbb{C},1}}:W_{1}(U_{\mathbb{C},d,n})\cong D(U_{\mathbb{C},d,n})\rightarrow G_{\mathbb{C}}(d,n)$. For two vector bundles $V_{1}$ and $V_{2}$ over the same base,
@@
$\mathbf{MTU}(d,1)_{2n}\cong Th(U_{\mathbb{C},d,n}\oplus U^{\perp}_{\mathbb{C},d,n})$
Since, $U_{\mathbb{C},d,n}\oplus U^{\perp}_{\mathbb{C},d,n}$ is a trivial $d+n$ dimensional complex bundle,
$\mathbf{MTU}(d,1)_{2n}\cong S^{2d+2n}\vee\Sigma^{2d+2n}G_{\mathbb{C}}(d,n)$
Letting $n\rightarrow\infty$ completes the proof. ∎
@@
This result follows by dualizing the corresponding result, Proposition 2.4, for the cohomology of $\mathbf{MTU}(d)$ from [9]. ∎
@@
There is an obstruction to lifting an element of $\pi_{2d}(\mathbf{MU})$ to $\pi_{2d}(\mathbf{MTU}(d-r))$ which lands in the homotopy group $\pi_{2d}(\mathbf{MTU}(d,r))$. In [9] we described the image of $\pi_{*}(\mathbf{MTU}(d))\otimes\mathbb{Q}\rightarrow\pi_{*}(\mathbf{MU})\otimes\mathbb{Q}$ as follows:
@@
Let $M^{2d}$ be a $d$-dimensional almost-complex manifold. The Chern characteristic classes $s_{\omega}(M^{2d})=0$ for all $\omega$ of length greater than $d-r$ if and only if the cobordism class $[M^{2d}]$ lifts to $\pi_{2d}(\mathbf{MTU}(d-r))\otimes\mathbb{Q}$. Geometrically, this implies there exists integer $c$ such that the cobordism class $c[M^{2d}]$ contains a manifold $N^{2d}$ with $r$ complex sections on $TN$.
We will now describe the image of the map $\pi_{*}(\mathbf{MTU}(d))\rightarrow\pi_{*}(\mathbf{MU})$. As stated in the introduction, we will call the torsion in the cokernel, the torsion obstructions. Thus we determine which cobordism classes of manifolds can be equipped with complex sections. To do so, we will study the $p$-local Adams-Novikov spectral sequence for $\mathbf{MTU}(d)$.
@@ 3. Computing the torsion obstruction
This section will be devoted to identifying which elements of $\pi_{*}(\mathbf{MU})$ can be lifted to $\pi_{*}(\mathbf{MTU}(d))$. We will use the following notational conventions. Then let $B_{i}\in MU_{*}(\mathbf{MU})\cong\pi_{*}(\mathbf{MU})\otimes H_{*}(\mathbf{MU})$, be the generators (as a $\pi_{*}(\mathbf{MU})$ algebra). These $B_{i}$ correspond to the elements $b_{i}\in H_{*}(\mathbf{MU})\cong\mathbb{Z}[b_{1},b_{2},...]$. There is an injective Hurewicz homomorphism $\pi_{*}(\mathbf{MU})\rightarrow H_{*}(\mathbf{MU})$. By slight abuse of notation, we call the elements of $\pi_{*}(\mathbf{MU})$ by their image in $H_{*}(\mathbf{MU})$. We begin by stating the following results about $MU$ theory originally developed by [10] and described in detail in [11] Appendix 2.1.
@@
  • A right unit, $\eta_{R}:\pi_{*}(\mathbf{MU})\rightarrow MU_{*}(\mathbf{MU})$, corresponding to the Hurewicz homomorphism, defined by the formula:
@@
For each prime $p$, there is an associative ring spectrum $\mathbf{BP}$ which is a retract of $\mathbf{MU}_{(p)}$, moreover, $\mathbf{MU}_{(p)}$ splits as copies of $\mathbf{BP}$. The homotopy group $\pi_{*}(\mathbf{BP})$ is $\mathbb{Z}_{(p)}[v_{1},v_{2},...]$ where $v_{i}\in\pi_{2(p^{i}-1)}(\mathbf{BP})$. The pair $(\pi_{*}(\mathbf{BP}),BP_{*}(\mathbf{BP}))$ is a Hopf algebroid. We will write $BP_{*}(\mathbf{BP})=\pi_{*}(\mathbf{BP})[t_{1},t_{2},...]$ with element $t_{i}$ in degree $2(p^{i}-1)$. The algebroid has the structure:
@@
  • A right unit, $\eta_{R}:\pi_{*}(\mathbf{BP})\rightarrow BP_{*}(\mathbf{BP})$, corresponding to the Hurewicz homomorphism, defined by the formula:
@@
The basis $m_{i}$ comes from $[\mathbb{C}P^{p^{i}-1}]/p$. The generators $v_{i}$ are defined by
@@
Let $X$ be a spectrum with torsion-free homology concentrated in even degrees. Let $E$ be either $\mathbf{MU}$ or $\mathbf{BP}$ for some prime $p$. Then the generalized theory $E_{*}(X)\cong\pi_{*}(E)\otimes_{\mathbb{Z}}H_{*}(X;\mathbb{Z})$ and is torsion-free concentrated in even degrees.
@@
Let $D$ be either $\mathbf{MU}$ or $\mathbf{BP}$ for some prime $p$. The map
$D_{*}(\mathbf{MTU}(d))\rightarrow D_{*}(\mathbf{MU})=\pi_{*}(E)[B_{1},B_{2},...]$
is an injection, which is an isomorphism in degrees less than $22d+$ 2. Moreover, both modules are concentrated in even degrees. The image of $D_{*}(\mathbf{MTU}(d))$ maps onto polynomials of degree $d$.
$D_{*}(\mathbf{MU})\rightarrow D_{*}(\overline{\mathbf{MTU}}(d))$
is the quotient map killing all polynomials of degree $d$ or less. These maps induce natural structures as $D_{*}(D)$ comodules.
We will use the Adams-Novikov spectral sequence as our main tool to compute the cokernel of the map $\pi_{*}(\mathbf{MTU}(d))\rightarrow\pi_{*}(\mathbf{MU})$. The $p$-local Adams-Novikov spectral sequence computes the homotopy groups of a finite type spectrum $X$ as follows, see [10] or [11].
@@
Moreover, this construction is natural with respect to maps of spectra. For an element in $E_{q}^{s,t}$, we will refer to $s$ as the resolution of the element and $t$ as the degree. The element, if it survives, corresponds to an element of $\pi_{t-s}$. Consider the Adams-Novikov spectral sequence for $\mathbf{MU}_{(p)}$. Since $\mathbf{MU}_{(p)}$ splits as a copy of $\mathbf{BP}$’s, $E_{2}^{s,*}(\mathbf{MU}_{(p)})=0$ for $s>0$. So the spectral sequence collapses, and $E_{2}^{0,*}(\mathbf{MU}_{(p)})\cong\pi_{*}(\mathbf{MU}_{(p)})$. In particular, it is a free $\mathbb{Z}_{(p)}$ module. Moreover, by basic properties of $\mathbf{Ext}$ groups,
$E_{2}^{0,*}(\mathbf{MTU}(d)_{(p)})\cong\mathbf{Hom}(BP_{*}(\mathbf{MTU}(d)_{(p)}),\pi_{*}(\mathbf{BP}))$
The image of $\pi_{*}(\mathbf{MTU}(d)_{(p)})\rightarrow\pi_{*}(\mathbf{MU}_{(p)})$ is concentrated in the 0 resolution. Thus, we first need to know the image of the map $E_{2}^{0,*}(\mathbf{MTU}(d))\rightarrow E_{2}^{0,*}(\mathbf{MU})$. This is the map
$\mathbf{Hom}_{BP_{*}(\mathbf{BP})}(\pi_{*}(\mathbf{BP}),BP_{*}(\mathbf{MTU}(d)_{(p)}))\rightarrow\mathbf{Hom}_{BP_{*}(\mathbf{BP})}(\pi_{*}(\mathbf{BP}),BP_{*}(\mathbf{MU}_{(p)}))$
This map has cokernel which is a free $\mathbb{Z}_{(p)}$ module. Since all differentials are zero in the Adams-Novikov spectral sequence for $\mathbf{MU}_{(p)}$, the following proposition is concluded:
@@
Moreover, $[M]$ lifts to $\pi_{*}(\mathbf{MTU}(d))$ if and only if there is no Adams differential originating from $x$.
Proof.
Let $[M]\in\pi_{*}(\mathbf{MU})$, where $p[M]\in\pi_{*}(\mathbf{MU})$ lifts to $\pi_{*}(\mathbf{MTU}(d))$. We may view $[M]$ as an element $z\in E_{2}^{0,*}(\mathbf{MU})$. We know $p[M]$ can be viewed as an element in $\pi_{*}(\mathbf{MTU}(d))$, which lifts to $y\in E_{2}^{0,*}(\mathbf{MTU}(d))$. By naturality of the Adams Novikov spectral sequence $y\mapsto pz$. Since the cokernel of $E_{2}^{0,*}(\mathbf{MTU}(d))\rightarrow E_{2}^{0,*}(\mathbf{MU})$ is free and the map is injective, there must be a unique element $x\mapsto z$. The element $x$ survives to the $E_{\infty}$ page if and only if all Adams differentials vanish on $x$. ∎
Thus, the differentials in the Adams-Novikov spectral sequence for $\mathbf{MTU}(d)$ correspond with torsion obstructions. We will use the cobar complex in order to compute the $E_{2}$ page. See Appendix 1 in [11] for more details. If we wished to extend this computation further, it would be beneficial to use a more sophisticated method. Recall that $BP_{*}(\mathbf{BP})$ has an action of $\pi_{*}(\mathbf{BP})$ given by multiplication by the image of $\eta_{R}$. Let $\overline{\Gamma}$ be the quotient of $BP_{*}(\mathbf{BP})$ the image of $\eta_{L}$.
@@
Let $(M,\psi)$ be a left $BP_{*}(\mathbf{BP})$ comodule. Define the cobar complex as the tensor product $C_{BP}^{s}(M)=\overline{\Gamma}^{\otimes s}\otimes_{\pi_{*}(\mathbf{BP})}M$. The differential in the complex is $d_{1}:C_{BP}^{s}(M)\rightarrow C_{BP}^{s+1}(M)$, given by:
@@
It will typically be convenient to view the cobar complex as the $E_{1}$ page of the Adams-Novikov spectral sequence where the differential corresponds to $d_{1}$. We observe that the 00 line of the $E_{2}$ page of the Adams-Novikov spectral sequence is generated by primitives, i.e. elements of the form $\psi(x)=1\otimes x$ where $\psi$ is the coaction. The following proposition will be used to compute $E_{2}(\mathbf{MTU}(d))$